• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 00:02:51 on April 8,2011

    Entry number 354720

    keyword=Swing shift summary

    Swing shift summary:

    Shift workers: J. P. Chen (SL), Nuruzzaman (TO), M. Shabestari(3rd)

    16:00 Beam study, hall in controlled access.
    17:00 Hall back to beam permit.
    20:00 Moved target to EMPTY to allow beam tuning
    21:00 Completed tuning beam, asked for beam position as in the white board. MCC informed us that they need to tune for Hall C
    21:30 MCC came back to us. Beam position tuning.
    21:40 Beam position good. Move target to BeO.
    22:05 Raster check with BeO
    22:10 Move to 4He, raster check.
    22:20 DAQ problem, TS11 (needed reboot from hareboot30), then different ROCs.
    23:00 Raster check, spot slightly larger than 2x2. Do another run, DAQ problem again.
    23:10 beam off.
    23:20 DAQ problem solved. wait for beam.
    23:35 beam back, raster check.
    23:45 beam off.
    23:50 beam back.
    23:55 start production run 3347.
    24:00 run 3347 still on-going.